# Alexandra Ciobotaru

### v2

@@ Abstract
Psychological corpora in NLP are collections of texts used to analyze human psychology, emotions, and mental health. These texts allow researchers to study psychological constructs, detect mental health issues and analyze emotional language. However, mental health data can be difficult to collect correctly from social media, due to suppositions made by the collectors. A more pragmatic strategy involves gathering data through open-ended questions and then assessing this information with self-report screening surveys. This method was employed successfully for English, a language with a lot of psychological NLP resources. However, this cannot be stated for Romanian, which currently has no open-source mental health corpus. To address this gap, we have created the first corpus for depression and anxiety in Romanian, by utilizing a form with 6 open-ended questions along with the standardized PHQ-9 and GAD-7 screening questionnaires. Consisting of the texts of 205 respondents and although it may seem small, PsihoRo is a first step towards understanding and analyzing texts regarding the mental health of the Romanian population. We employ statistical analysis, text analysis using Romanian LIWC, emotion detection and topic modeling to show what are the most important features of this newly introduced resource to the NLP community.
@@ 2 Related Work
[35] created open-ended personality questions to explore how natural language processing can aid in predicting neuroticism from people’s responses. They gathered over 30,000 sentences written by 425 adults in Korean and used KoBERT to see which types of questions best revealed neuroticism and related traits like depression and dependency.
[34] created the multi-modal Real World Worry Waves Dataset (RW3D), in English, which tracks the psychological effects of the COVID-19 pandemic in the UK over three consecutive years, 2020, 2021 and 2022, through repeated surveys containing also free-text responses.
[10] used open-ended interview responses from more than 58,000 job candidates along with self-assessed personality (HEXACO) and employed “InterviewBERT” to deduce personality traits from textual replies. This study was also done in English.
[30] used open-ended questionnaires to study young and middle aged adults’ responses regarding their own mental health, by employing NLP computational methods. They analyzed their 1106 responses and found various differences in semantics on the two study groups.
[3] combined traditional linguistic analysis (LIWC and sentiment analysis) with NLP and machine learning to find a relationship between text style and depression, anxiety, and stress during the COVID-19 pandemic. They analyzed written narratives from 883 participants combined with their DASS-21 evaluation scores and found a connection between increased symptom severity and the greater use of first-person pronouns and negative emotion words.
[19] collected responses in English from 411 participants who answered word-response questions related to their mental health and completed also the GAD-7 and PHQ-9 rating scales. Their analyses conclude that individuals’ language reflects their mental health status, and combining verbal responses with quantitative measures can enhance the comprehension and assessment the risk of depressive and anxiety symptoms in both clinical practice and research contexts.
In terms of combing emotion detection and LIWC term analysis to mental health texts, there are several studies which employ such analysis on the English language: [26], [36], [2].
There are several datasets created by combining gathering text through open ended questionnaires together with a screening tool for detecting the risk of a potential mental problem, but mainly centered on the English language. For Romanian, currently there are no open source psychological corpora available, and this fact was the main driving force behind our study, along with the curiosity of applying emotional analysis on such a corpus.
@@ 3 Data Collection
By using this survey collection methodology, participants provide their informed consent for the collection of data, as opposed to social media collected data, where data is collected without asking the participants for their consent [4].
The data collection methodology consists of a completely anonymous survey, organized in three parts. The first part of the survey consists of six open-ended questions. Its scope is to accumulate as much written text as possible. Out of these six questions, three of them touch positive topics and three of them negative topics. The survey was written in Romanian for Romanian speakers only. Below are English translations for the six open-ended questions:
@@
    To what extent do you think social media affects your emotional state, especially when you see negative news? "În ce măsură crezi că rețelele sociale îți afectează starea emoțională, mai ales când vezi știri negative?" (negative)
@@
    What is an important lesson about resilience and adaptation that you have learned recently? "Care este o lecție importantă despre reziliență și adaptare pe care ai învățat-o recent?" (positive)
@@
    How much do you think politics affects people’s mental health? Can you describe a time when you felt stressed because of political decisions? "Cât de mult crezi că politica afectează sănătatea mentală a oamenilor? Poți să detaliezi un moment în care ai simțit stres din cauza deciziilor politice?" (negative)
@@
    What role do you think community or social support plays in maintaining mental health during difficult times? "Ce rol crezi că are comunitatea sau sprijinul social în menținerea sănătății mentale în vremuri dificile? "(positive)
@@
    How were you affected by discussions about war, the economic crisis, elections, or other tense topics? "Cum ai fost afectat de discuțiile despre război, criza economică, alegeri sau alte subiecte tensionate?" (negative)
@@
    How do you recharge your batteries after a stressful period? "Cum îți reîncarci bateriile după o perioadă stresantă?" (positive)
We wanted to create a corpus of text data written by persons prone to suffering from depression and/or anxiety. Usually psychological corpuses for NLP are usually gathered from social media through searching searching different keywords in social media posts that could indicate depression and/or anxiety. By discussing with medical professionals in the psychiatric field we concluded that this type of gathering psychological data is speculative. Thus, we added to our survey two more parts, one that assesses if the respondent is prone to depression by incorporating PHQ-9 [21] questions, and another part which assesses if the respondent is prone to anxiety by incorporating GAD-7 [32] questions. Both GAD-7 and PHQ-9 contain questions (7 for GAD-7 and 9 for PHQ-9) which ask how often the respondent has been bothered by certain problems, and the response options correspond to a numerical score: Not at all = 0, Several days = 1, More than half the days = 2, Nearly every day = 3.
The severity cut-offs for depression based on PHQ-9 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe, while the severity cut-offs for anxiety based on GAD-7 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe.
The resulting questionnaire is completely anonymous, no personal data is requested or stored (such as name, surname, email address, IP). We did not collect demographic data for the same reason: to gain the trust of respondents and also to keep the total time of form completion short.
The resulting data consists of 205 responses, collected over an eight-month period from March 2025 to October 2025. This time frame includes two major stressful events for Romanians, presidential elections and the tax increase burden generated by the country’s new leadership.
@@ 4 Methodology
  • Depression Risk Group: 37 respondents (18%) scored higher than 10 on PHQ-9;
  • Anxiety Risk Group: 36 respondents (17.56%) scored higher than 10 on GAD-7;
  • No Risk Group: 28 respondents (13.66%) scored 0 for at least one out of the two questionnaires.
@@
In parallel to these anlyses we have created an emotion detection model by fine-tuning the Romanian BERT [12] on the REDv2 dataset [8], for multi-label text classification, using the Huggingface transformers pipeline [^1] https://huggingface.co/docs/transformers. We used this model to predict the emotions in PsihoRo, thus gathering emotional information for the text written by each population group mentioned above. The resulting model had an F1 score of 66.85% computed on the REDv2 test set.
@@
In order to gain interpretable results, the words resulted from the most important topics were further expanded into their main reformulated ideas using GPT-5.
@@ 5.1 PsihoRo Descriptive Statistics
By analyzing the length of the answers, we questioned whether people with risk of depression and/or anxiety tend to write less. The answer to this question is no, because the computed T-statistic was 1.55 between the depression risk population and the no risk population, and 2.26 between the anxiety risk population and the no risk population, meaning that no significant difference was found.
In Figure 1 we show the distribution plots of both PHQ-9 and GAD-7 scores. The data showed that the distributions are skewed to the left, with the majority of observations concentrated at lower values and a longer tail extending toward the higher end of the scale, meaning that most respondents have minimal and mild risk of depression and anxiety.
@@ 5.2 LIWC
Surprisingly, the use of negations is negatively correlated with both PHQ-9 and GAD-7 scores, suggesting that participants with higher depression and anxiety scores tend to use fewer negations in their responses. Individuals who score higher on these assessments tend to use more words in the Tentative category, which indicates uncertainty, supported by the correlation coefficients of .204** for PHQ-9 and .213** for GAD-7. Comparatives also show a positive correlation with the depression score (.167*), indicating comparative thinking, which is often linked to depressive rumination. However, another marker for rumination, the high frequency of Past focus verbs, is not observed in our data. There is a positive correlation between the use of Body-related words and PHQ-9 (200**) and GAD-7 (180**) scores, indicating that participants with higher scores frequently use more language related to bodily sensations. This may reflect increased somatic awareness, a common trait in depression and anxiety [25]. Participants with higher PHQ-9 and GAD-7 scores discuss Achievement less frequently (-.163* and- .192**, respectively), which could indicate low motivation. In addition, the negative correlation with Leisure-related words (-.145* and -.169*, respectively) may stem from anhedonia, a common symptom of depression. Words in the Death category show a significant positive correlation with depression (.138*), but not with anxiety (0.029), indicating a greater preoccupation with mortality in participants with higher depression scores. In contrast, the Family category is positively correlated with anxiety scores (r = 0.145*), suggesting that individuals with higher anxiety are more focused on family-related issues.
@@ 5.3 Emotion Analysis
Surprise is an emotion that is completely absent from the PsihoRo data. The following emotions are more common, in sequence of appearance: anger, joy, fear, sadness, trust, and the most common, neutral. The emotional pattern for the more pessimistic questions is displayed in the center, while the emotional pattern for the more optimistic questions is displayed on the right. Emotions such as sadness and anger are on the rise on the left, while joy and trust are more prevalent on the right.
The emotions present in the texts written by the three primary demographic groups mentioned at the start of Section 4 were also examined in Figure 6. It can be seen that besides the neutral, which is strongly present across all groups
@@
We also tried to create several regression models, based on the textual PsihoRo data gathered and the GAD-7 and PHQ-9 scores, using the Romanian BERT for text vectorization and also adding emotional information as features, but all our experiments had negative R2.
@@ 5.4 Topic Modeling
We applied the topic modeling analysis described in Section 4 for each of the three control groups studied in this article. The main topics are derived from the six questions asked the form. However, some differences appear between the three groups of respondents, which can be found in the following subsections [^2] The following subsections were generated using GPT-5, based on the words generated by the topic modeling operation, some of which present in text in bold..
@@ 5.4.1 No Risk Group
Topic 1. This theme highlights how individuals perceive the current political and social period in a critical and sometimes negative way. The words “negative”, “decisions”, “politics”, “measure”, “news” and “period” indicate a concern for the effects of government decisions and how they are reflected in the media. The presence of the term “sport” suggests a possible comparison between different areas of public life.
Topic 2. The theme reflects the complex interaction between information conveyed through news, political life and how they influence the general state of people. Terms such as “politics”, “news”, “state”, “people” and “politics” indicate an increased attention to how media discourse shapes collective perceptions and emotions.
Topic 3. This topic addresses how individuals interpret and process information related to elections and political decisions. The terms “measure”, “depends”, “people”, “time”, “elections” and “example” suggest a personal analysis of the electoral context and an attempt to understand the relationship between information, time and decision.
Topic 4. The theme highlights the importance of authentic human interaction and the balance between the inner and social dimensions of the individual. The words "listen", "outer", "inner", "community" and "matter" suggest an openness to empathetic communication and a valorization of social belonging.
Topic 5. This topic focuses on the emotional dimension of everyday experiences, indicating how individuals manage their affective states and reactions in common situations. Words such as “affect”, “try”, “day”, “go” and “enough” express a constant effort to adapt and maintain psychological balance in the face of daily pressures by self regulation and personal reflection.
@@ 5.4.2 Depression Risk Group
Topic 1. This theme captures the perception of fatigue, daily routine and the constant effort that individuals put into their personal and professional lives. Associated terms such as “sleep”, “hard”, “walking” and “people” indicate a general state of exhaustion and the need for balance between personal and social life.
Topic 2. This theme highlights the importance of psychological resilience and the ability to adapt to stressful or unpredictable situations. Words such as “resilience”, “situation”, “time”, “community” and “important” suggest a solution-oriented approach and maintaining a state of balance in contexts of uncertainty.
Topic 3. The topic focuses on how individuals relate to political processes and the electoral climate. The terms “politics”, “elections”, “people”, “moment” and “listening” indicate both interest in political events and their influence on social attitudes and conversations.
Topic 4. This topic captures the potentially harmful impact of social networks on interpersonal relationships and community cohesion. The presence of the words “negative”, “social”, “networks”, “community” and “socialization” underlines the perception of a degradation of authentic communication and well-being caused by the digital environment.
Topic 5. This theme reflects the concern for affective states and how social and economic factors contribute to the intensification of collective anxiety. The terms “feel”, “affect”, “cause”, “anxiety”, “country” and “family” suggest an emotional perspective on everyday experiences and the general social context.
@@ 5.4.3 Anxiety Risk Group
Topic 1. This topic captures how major events, such as war, economic crises or socio-political changes, affect the perception of daily life. Words such as “affect”, “crisis”, “war”, “elections”, “life” and “time” indicate the concern for external instability and its influence on personal balance.
Topic 2. This topic highlights the close connection between political factors and individual emotional state. Terms such as “politics”, “stress”, “affect”, “feel”, “cause” and “family” suggest a direct relationship between socio-political events and the level of anxiety or psychological tension felt.
Topic 3. This topic addresses the role of community and collective resilience in the face of the constant flow of negative information from the media. The words “resilience”, “news”, “negative”, “community”, “important” and “trying” indicate an attempt to find the balance between information and protecting one’s mental state.
Topic 4. This topic explores the effects of social media on psychological state, with a focus on anxiety and social pressure. Terms such as “social”, “anxiety”, “state”, “hard”, “networks” and “present” suggest a connection between the intensive use of social networks and the degradation of emotional state. The topic reflects how virtual interactions can accentuate psychological discomfort and feelings of isolation.
Topic 5. This theme emphasizes the connection between people, community and the sense of belonging. The words “people”, “community”, “listen”, “dear” and “happen” highlight the importance of close relationships and solidarity in maintaining emotional balance.

### v3

@@ Abstract
Psychological corpora in NLP are collections of texts used to analyze human psychology, emotions, and mental health. These texts allow researchers to study psychological constructs, detect mental health issues and analyze emotional language. However, mental health data can be difficult to collect correctly from social media, due to suppositions made by the collectors. A more pragmatic strategy involves gathering data through open-ended questions and then assessing this information with self-report screening surveys. This method was employed successfully for English, a language with a lot of psychological NLP resources. However, this cannot be stated for Romanian, which currently has no open-source mental health corpus. To address this gap, we have created the first open-source corpus for depression and anxiety in Romanian, by utilizing a form with 6 open-ended questions along with the standardized PHQ-9 and GAD-7 screening questionnaires. Consisting of the texts of 205 respondents and although it may seem small, PsihoRo is a first step towards understanding and analyzing texts regarding the mental health of the Romanian population. We employ statistical analysis, text analysis using Romanian LIWC, emotion detection and topic modeling to show what are the most important features of this newly introduced resource to the NLP community. The data is publicly available at https://huggingface.co/datasets/Alegzandra/PsihoRo.
@@ 2 Related Work
Previous research has gathered open-ended responses from participants to explore how expressions of mental health influence language. [39] created open-ended personality questions to explore how natural language processing can aid in predicting neuroticism from people’s responses. They gathered over 30,000 sentences written by 425 adults in Korean and used KoBERT to see which types of questions best revealed neuroticism and related traits like depression and dependency. [37] created the multi-modal Real World Worry Waves Dataset (RW3D), in English, which tracks the psychological effects of the COVID-19 pandemic in the UK over three consecutive years, 2020, 2021 and 2022, through repeated surveys containing also free-text responses. [11] used open-ended interview responses from more than 58,000 job candidates along with self-assessed personality (HEXACO) and employed “InterviewBERT” to deduce personality traits from textual replies. This study was also done in English. [33] used open-ended questionnaires to study young and middle aged adults’ responses regarding their own mental health, by employing NLP computational methods. They analyzed their 1106 responses and found various differences in semantics on the two study groups.
In terms of combining emotion detection and LIWC term analysis in mental health texts, several studies employ such analysis in English: [29], [40], [2]. [4] combined LIWC features [6], sentiment analysis and NLP to explore the relationship between language and mental health issues such as depression, anxiety, and stress during the COVID-19 pandemic. They analyzed written narratives from 883 participants, alongside their DASS-21 scores, and found an association between increased symptom severity and greater use of first-person pronouns and negative emotion words. [21] collected responses in English from 411 participants who answered word-response questions related to their mental health and completed the GAD-7 and PHQ-9 rating scales. Their analyses show that individuals’ language reflects their mental health status, and combining verbal responses with quantitative measures can enhance the comprehension and assessment of the risk of depressive and anxiety symptoms in both clinical practice and research contexts.
While several datasets have been created using text gathered from open-ended questionnaires, combined with screening tools to assess the risk of mental health issues, most of these resources are focused on the English language. Currently, there are no open-source psychological corpora available in Romanian, and this was the main driving force behind our study, along with our interest in applying emotional analysis to such a corpus.
@@ 3 Data Collection
As opposed to social media collected data, which is collected without participants’ consent, we collected data from participants who provided their informed consent for data collection [5]. The data collection methodology consists of a completely anonymous survey, organized in three parts. The first part of the survey consists of six open-ended questions, designed to gather extensive written responses. Out of these six questions, three touch on positive topics and three on negative topics. The survey was written in Romanian for Romanian speakers only. We further present the six open-ended questions, in both English and Romanian:
@@
    To what extent do you think social media affects your emotional state, especially when you see negative news? “În ce măsură crezi că rețelele sociale îți afectează starea emoțională, mai ales când vezi știri negative?” (negative)
@@
    What is an important lesson about resilience and adaptation that you have learned recently? “Care este o lecție importantă despre reziliență și adaptare pe care ai învățat-o recent?” (positive)
@@
    How much do you think politics affects people’s mental health? Can you describe a time when you felt stressed because of political decisions? “Cât de mult crezi că politica afectează sănătatea mentală a oamenilor? Poți să detaliezi un moment în care ai simțit stres din cauza deciziilor politice?” (negative)
@@
    What role do you think community or social support plays in maintaining mental health during difficult times? “Ce rol crezi că are comunitatea sau sprijinul social în menținerea sănătății mentale în vremuri dificile?” (positive)
@@
    How were you affected by discussions about war, the economic crisis, elections, or other tense topics? “Cum ai fost afectat de discuțiile despre război, criza economică, alegeri sau alte subiecte tensionate?” (negative)
@@
    How do you recharge your batteries after a stressful period? “Cum îți reîncarci bateriile după o perioadă stresantă?” (positive)
Usually, psychological corpora for NLP are gathered from social media through searching for different keywords in online posts that could indicate depression and anxiety. However, after consulting with medical professionals in the psychiatric field, we have concluded that this method of collecting psychological data may not be entirely accurate. Thus, we added two more parts to our survey: one assessing respondents’ risk of depression using the PHQ-9 [23], and another assessing their risk of anxiety using the GAD-7 [35]. Both GAD-7 and PHQ-9 contain questions (7 for GAD-7 and 9 for PHQ-9) that ask the participants how often they have experienced specific problems, and the response options correspond to a numerical score: Not at all = 0, Several days = 1, More than half the days = 2, Nearly every day = 3. The severity cut-offs for depression based on PHQ-9 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe, while the severity cut-offs for anxiety based on GAD-7 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe.
The resulting questionnaire is completely anonymous, and no personal data is requested or stored (such as name, surname, email address, or IP). We did not collect demographic data for the same reason: to build trust with respondents and keep the time required to complete the form short.
The resulting data consists of 205 responses, collected over an eight-month period from March 2025 to October 2025. This time frame includes two major stressful events for Romanians - presidential elections and the tax increase burden generated by the country’s new leadership.
@@ 4 Methodology
  • Depression Risk Group: 37 respondents scored higher than 10 on PHQ-9 (18%);
  • Anxiety Risk Group: 36 respondents scored higher than 10 on GAD-7 (17.56%);
  • No Risk Group: 28 respondents scored 0 for at least one out of the two questionnaires (13.66%).
@@
To complement these analyses, we have created an emotion detection model by fine-tuning the Romanian BERT [13] on the REDv2 dataset [9], for multi-label text classification, using the Huggingface transformers pipeline [^1] https://huggingface.co/docs/transformers. We used this model to predict the emotions in PsihoRo, thus gathering emotional information for the text written by each population group mentioned above. The resulting model had an F1 score of 66.85% computed on the REDv2 test set.
@@ 5.1 PsihoRo Descriptive Statistics
We analyzed the length of responses to determine whether individuals at risk of depression or anxiety tend to write less. Our findings indicate that this is not the case for Romanian. The computed T-statistic was 1.55 when comparing the depression risk group with the group not at risk, and 2.26 when comparing the anxiety risk group with the group not at risk. These results suggest that there is no significant difference in response length between these populations.
In Figure 1 we show the distribution plots of both PHQ-9 and GAD-7 scores. The data indicate that the distributions are skewed to the left, with the majority of observations concentrated at lower values and a longer tail extending toward the higher end of the scale, meaning that most respondents have minimal and mild risk of depression and anxiety.
@@ 5.2 LIWC
Surprisingly, the use of negations is negatively correlated with both PHQ-9 and GAD-7 scores, suggesting that participants with higher depression and anxiety scores tend to use fewer negations in their responses. Individuals who score higher on these assessments tend to use more words in the Tentative category, which indicates uncertainty, supported by the correlation coefficients of .204** for PHQ-9 and .213** for GAD-7. Comparatives also show a positive correlation with the depression score (.167*), indicating comparative thinking, which is often linked to depressive rumination. However, another marker for rumination, the high frequency of Past focus verbs, is not observed in our data. There is a positive correlation between the use of Body-related words and PHQ-9 (200**) and GAD-7 (180**) scores, indicating that participants with higher scores frequently use more language related to bodily sensations. This may reflect increased somatic awareness, a common trait in depression and anxiety [28]. Participants with higher PHQ-9 and GAD-7 scores discuss Achievement less frequently (-.163* and- .192**, respectively), which could indicate low motivation. In addition, the negative correlation with Leisure-related words (-.145* and -.169*, respectively) may stem from anhedonia, a common symptom of depression ([15]). Words in the Death category show a significant positive correlation with depression (.138*), but not with anxiety (0.029), indicating a greater preoccupation with mortality in participants with higher depression scores. In contrast, the Family category is positively correlated with anxiety scores (r = 0.145*), suggesting that individuals with higher anxiety are more focused on family-related issues.
@@ 5.3 Emotion Analysis
Surprise is an emotion that is completely absent from the PsihoRo data. The following emotions are more common, in order of appearance: anger, joy, fear, sadness, trust, and the most common, neutral. The emotional pattern for the more pessimistic questions is displayed in the center of Figure 5, while the emotional pattern for the more optimistic questions is displayed on the right. Emotions such as sadness and anger are on the rise on the Pessimistic Questions, while joy and trust are more prevalent on the Optimistic Questions.
The emotions present in the texts written by the three primary demographic groups mentioned at the beginning of Section 4 are presented in Figure 6. While the neutral emotion is strongly represented across all groups, the population at risk for depression shows a high prevalence of sadness. In addition, the group at risk for anxiety exhibits emotions of sadness, fear and trust.
@@ 5.4 Topic Modeling
We applied the topic modeling analysis described in Section 4 for each of the three population groups studied in this work. The main topics are derived from the six open questions asked in the form. However, some differences appear among the three groups of respondents, as discussed in the following subsections.
@@ 5.4.1 No Risk Group
Topic 1 highlights how individuals perceive the current political and social period in a critical and sometimes negative way. The words “negative”, “decisions”, “politics”, “measure”, “news” and “period” indicate a concern for the effects of government decisions and how they are reflected in the media. The presence of the term “sport” suggests a possible comparison between different areas of public life.
Topic 2 reflects the complex interaction between information conveyed through news, political life, and how they influence the general state of the population. Terms such as “politics”, “news”, “state”, “people” and “politics” indicate an increased attention to how media discourse shapes collective perceptions and emotions.
Topic 3 addresses how individuals interpret and process information related to elections and political decisions. The terms “measure”, “depends”, “people”, “time”, “elections” and “example” suggest a personal analysis of the electoral context and an attempt to understand the relationship between information, time and decision.
Topic 4 highlights the importance of authentic human interaction and the balance between the inner and social dimensions of the individual. The words “listen”, “outer”, “inner”, “community” and “matter” suggest an openness to empathetic communication and highlight the value of social belonging.
Topic 5 focuses on the emotional dimension of everyday experiences, indicating how individuals manage their affective states and reactions in common situations. Words such as “affect”, “try”, “day”, “go” and “enough” express a constant effort to adapt and maintain psychological balance in the face of daily pressures by self-regulation and personal reflection.
@@ 5.4.2 Depression Risk Group
Topic 1 captures the perception of fatigue, daily routine and the constant effort that individuals put into their personal and professional lives. Associated terms such as “sleep”, “hard”, “walking” and “people” indicate a general state of exhaustion and the need for balance between personal and social life.
Topic 2 highlights the importance of psychological resilience and the ability to adapt to stressful or unpredictable situations. Words such as “resilience”, “situation”, “time”, “community” and “important” suggest a solution-oriented approach and maintaining a state of balance in contexts of uncertainty.
Topic 3 focuses on how individuals relate to political processes and the electoral climate. The terms “politics”, “elections”, “people”, “moment” and “listening” indicate both interest in political events and their influence on social attitudes and conversations.
Topic 4 captures the potentially harmful impact of social networks on interpersonal relationships and community cohesion. The presence of the words “negative”, “social”, “networks”, “community” and “socialization” underlines the perception of a degradation of authentic communication and well-being caused by the digital environment.
Topic 5 reflects the concern for affective states and how social and economic factors contribute to the intensification of collective anxiety. The terms “feel”, “affect”, “cause”, “anxiety”, “country” and “family” suggest an emotional perspective on everyday experiences and the general social context.
@@ 5.4.3 Anxiety Risk Group
Topic 1 captures how major events, such as war, economic crises or socio-political changes, affect the perception of daily life. Words such as “affect”, “crisis”, “war”, “elections”, “life” and “time” indicate the concern for external instability and its influence on personal balance.
Topic 2 highlights the close connection between political factors and individual emotional state. Terms such as “politics”, “stress”, “affect”, “feel”, “cause” and “family” suggest a direct relationship between socio-political events and the level of anxiety or psychological tension felt.
Topic 3 addresses the role of community and collective resilience in the face of the constant flow of negative information from the media. The words “resilience”, “news”, “negative”, “community”, “important” and “trying” indicate an attempt to find the balance between information and protecting one’s mental state.
Topic 4 explores the effects of social media on psychological state, with a focus on anxiety and social pressure. Terms such as “social”, “anxiety”, “state”, “hard”, “networks” and “present” suggest a connection between the intensive use of social networks and the degradation of emotional state. The topic reflects how virtual interactions can accentuate psychological discomfort.
Topic 5 emphasizes the connection between people, community and the sense of belonging. The words “people”, “community”, “listen”, “dear” and “happen” highlight the importance of close relationships and solidarity in maintaining emotional balance.